OpenStudy (anonymous):

What is the exact value of the expression the square . - the square root of 24. + the square root 54.root of 294? Simplify if possible

OpenStudy (anonymous):

@Koikkara yo i need help again \[\sqrt{294}-\sqrt{24}+\sqrt{54}\]

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@taylorlee pls help pls

OpenStudy (anonymous):

The easiest way to solve these problems are to factor each of those numbers out. When you factor them out you should find at least one number that multiplies by itself to get the original number. in other words the square. For example : the square root of 16 is 4 because 4 x 4 =16. If you didn't know this (i hope you did) then you could have factored out 16 or: 1,2,4,8,16. I would suggest that you do this for all of your numbers and then simplify your equation completely. Hope this helps!!!!